OpenStudy (anonymous):

integral of 1/(cos^2x tanx ) dx

OpenStudy (turingtest):

what else can we write 1/cos^2 as?

OpenStudy (turingtest):

please don't just give the answer whoever is typing

OpenStudy (anonymous):

\[\frac{1}{\cos ^{2}x \tan x}=\frac{1}{\cos x \sin x}\]

OpenStudy (turingtest):

^there is a better way

OpenStudy (anonymous):

=1/cos^2 x .tanx =1/(sin x . cos x) =cosecx.secx integration of this will be ln|tan x | am i right???

OpenStudy (anonymous):

\[u=\tan x\]

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

thats probably the best way

OpenStudy (anonymous):

@turingtest- what other way is there...

OpenStudy (turingtest):

\[{1\over\cos^2x\tan x}={\sec^2x\over\tan x}\]\[u=\tan x\]is the best way in my opinion

OpenStudy (turingtest):

how do you guys see\[\int \sec x\csc xdx\]as\[\ln|\tan x|+C\]anyway? cuz I don't see it

OpenStudy (anonymous):

neither do I know the derivation of that,i had an integral table given to me by one of the users here ,i looked upto to that hence it says that the integral will give ln|tan x|

OpenStudy (anonymous):

WolframAlph agrees with @TuringTest on the substitution.

OpenStudy (turingtest):

well I prefer integrals I can prove, and I don't know the proof of that I'm playing with it right now

OpenStudy (turingtest):

oh yeah, I promise this problem is probably "supposed" to be done by the sub I posted but there's more than one way to skin a cat!

OpenStudy (anonymous):

i agree...

OpenStudy (anonymous):

\[u = \tan x\]\[du = \sec^{2} xdx\]\[\int\limits \sec x \csc x dx=\int\limits \frac{1}{u} du\]

OpenStudy (anonymous):

Why do you want to skin a cat anyways?

OpenStudy (anonymous):

They are not that bad in general :P

OpenStudy (turingtest):

I don't know, satellite always says that, lol

OpenStudy (turingtest):

ok I see it now, but it seems a rather roundabout way to do it how would you know to sub u=tanx right there ? that way seems more counter-intuitive to me, but that's just my opinion

OpenStudy (anonymous):

Turing: If you have to ask, you'll never know. If you know, you need only ask. ;)

OpenStudy (anonymous):

true, the best way is sub u=tan x from the beginning

OpenStudy (turingtest):

ah-so sensei.... ;)